# MiniLLM: Knowledge Distillation of Large Language Models

### v5

@@ Abstract
Figure 1: The comparison of MiniLLM with the sequence-level KD (SeqKD; skd, alpaca, vicuna, ITGPT4, false_imitate, lima) in terms of the average GPT-4 feedback score on our evaluation sets. Left: GPT-2-1.5B as the teacher model and GPT-2 125M, 340M, 760M as the student models. Middle: GPT-J 6B as the teacher model and GPT-2 760M, 1.5B, GPT-Neo 2.7B as the student models. Right: OPT 13B as the teacher and OPT 1.3B, 2.7B, 6.7B as the student models.
@@ 1 Introduction
With the rapid development of large language models (LLMs; gpt3, plmsurvey, foundation_model, palm, gpt4), a common technique to reduce their high computational resource demand is knowledge distillation (KD; kd), where we train a small student model with supervision from a large teacher model. Two categories of KD are commonly applied: black-box KD, where only the teacher-generated texts are accessible, and white-box KD, where the teacher model’s output distribution or intermediate hidden states are also available [kdsurvey]. Recently, black-box KD has shown promising results in fine-tuning small models on the prompt-response pairs generated by LLM APIs [alpaca, vicuna, lamini, ITGPT4]. With the emergence of more open-source LLMs [opt, llama], white-box KD becomes more valuable for both research communities and industry sectors because student models receive better signals from the output distribution and hidden states of teacher models, thereby potentially resulting in higher performance. However, white-box KD approaches are mostly studied for small (<< 1B parameters) language understanding models [distilbert, minilm], while white-box KD for LLMs is yet to be explored.
In this work, we investigate white-box KD of LLMs where the output distribution of the teacher model is available. We argue that the standard KD objectives [skd, lightpaff, vicuna, alpaca] are sub-optimal for LLMs that perform tasks in a generative manner. Given the teacher distribution $p(\bm{y}|\bm{x})$ and the student distribution $q_{\theta}(\bm{y}|\bm{x})$ parameterized by $\theta$, standard KD objectives (including several variants for sequence-level models) essentially minimize the approximated forward Kullback-Leibler divergence (KLD) between the teacher and the student distribution, termed as $\operatorname{KL}[p||q_{\theta}]$, which forces $q_{\theta}$ to cover all modes of $p$. For text classification tasks, $\operatorname{KL}[p||q_{\theta}]$ works well because the output space usually consists of a finite number of classes such that both $p(\bm{y}|\bm{x})$ and $q_{\theta}(\bm{y}|\bm{x})$ have few modes. However, for open-ended text generation tasks, which is usually the case of LLM applications, the output spaces are much more complex and $p(\bm{y}|\bm{x})$ can contain many more modes than what $q_{\theta}(\bm{y}|\bm{x})$ can express due to the limited model capacity. Minimizing forward KLD causes $q_{\theta}$ to assign unreasonably high probabilities to the void regions of $p$ [malinin2019reverse] and produces very unlikely samples under $p$ during free-run generation [huszar2015not].
@@
To alleviate this problem, we propose to minimize reverse KLD, $\operatorname{KL}[q_{\theta}||p]$, widely used in computer vision [selfkd] and reinforcement learning [kd_policy_kd]. Compared to $\operatorname{KL}[p||q_{\theta}]$, minimizing $\operatorname{KL}[q_{\theta}||p]$ causes $q_{\theta}$ to seek the major modes of $p$, and assign low probabilities to $p$’s void regions [divergence_measure], as illustrated in Figure 2 and discussed in Section 2.1. In text generation, this means that the student avoids learning too many long-tail variants [top_p] in the teacher’s distribution to focuses on the generation correctness, which is critical in practical scenarios that require truthfulness and reliability [halu_survey]. To optimize $\min_{\theta}\operatorname{KL}[q_{\theta}||p]$, as shown in Section 2.2, we derive its gradient with Policy Gradient [policy_gradient] and adopt an on-policy training approach. To further stabilize and accelerate training, we propose (1) single-step decomposition to reduce variance, (2) teacher-mixed sampling to alleviate reward hacking, and (3) length normalization to eliminate the length bias. Finally, we introduce the overall KD algorithm in Section 2.3. Our student models are named MiniLLM, indicating our method is suitable for compressing large (generative) language models.
We apply our method to various generative language models [gpt2, opt, llama] with sizes ranging from 120M to 13B in the instruction-following setting [t0, flan] that covers a large range of NLP tasks. We use 5 datasets with Rouge-L [rouge], the GPT-4 feedback, and human judgment for evaluation. Experiments show that MiniLLM consistently outperforms standard KD baselines on all the datasets and scales up well from 120M to 13B models (see Figure 1). More analysis shows that MiniLLM yields lower exposure bias, better calibration, and higher long response generation performance, with neglectable loss of diversity.
@@ 2 Method
We consider conditional text generation where the model produces a response $\bm{y}=\{y_{t}\}_{t=1}^{T}$ conditioning on a prompt $\bm{x}$ sampled from the distribution $p_{\bm{x}}$, which is typically how LLMs perform tasks. We formulate KD as an optimization problem to minimize the difference between a fixed teacher model distribution $p(\bm{y}|\bm{x})$ and a student model distribution $q_{\theta}(\bm{y}|\bm{x})$ parameterized by $\theta$. The standard KD methods approximately [^1] We say “approximately” because for word-level KD, $\bm{y}$ is sampled from the real distribution, not the teacher distribution. For a strong enough teacher model, we can consider the two distributions approximately the same. minimize the forward KLD: $\operatorname{KL}[p||q_{\theta}]=\mathop{\mathbb{E}}_{\bm{x}\sim p_{\bm{x}},\bm{y}\sim p^{\prime}}\log\frac{p(\bm{y}|\bm{x})}{q_{\theta}(\bm{y}|\bm{x})}$, where $p^{\prime}$ can be real data distribution (word-level KD) or teacher distribution $p$ (sequence-level KD). Though widely used, $\operatorname{KL}[p||q_{\theta}]$ tends to overestimate the void regions of $p$ in text generation tasks when $q_{\theta}$ is insufficiently expressive [tailor]. KD for LLMs fits the case because LLMs perform tasks in a generative manner, such that the low-capacity student models cannot perfectly imitate the complex text generation distribution of the teacher models or humans.
@@ 2.1 MiniLLM: Knowledge Distillation with Reverse KLD
Minimizing reverse KLD has been shown to cause the mode-seeking behavior in generative modeling [huszar2015not, f-gan, chen2018symmetric, selfkd], where $q_{\theta}$ assigns high probabilities to $p$’s large modes and ignore the small ones (illustrated in a toy experiment in Figure 2). In this work, we first study this property for KD of LLMs in text generation. Minimizing forward KLD causes $q_{\theta}$ to place large probability masses on the zero-probability regions of $p$, corresponding to the generation of low-quality text in practice, while reverse KLD focuses on $p$’s major modes, which is crucial to ensure the correctness and faithfulness of text generation. As illustrated in Figure 3, unlike sequence-level KD that minimizes forward KLD [skd, alpaca], MiniLLM that minimizes reverse KLD does not force $q_{\theta}$ to fit all $\bm{y}$ sampled from the teacher distribution $p$. Instead, it encourages the student to generate samples preferred by the teacher within its own capacities, which is more possible to achieve. Interestingly, we also find another perspective of understanding MiniLLM motivated by Inverse Reinforcement Learning [maxent_irl]. We present the related derivation in Appendix A.1.
@@ 2.2 On-Policy Distillation
We notice that the gradient of the objective function $\mathcal{L}(\theta)$ in Equation (1) can be derived using the Policy Gradient Theorem [reinforce, max_ent_rl] for on-policy optimization:
@@
where $T=|\bm{y}|$ and $R_{t}=\sum_{t^{\prime}=t}^{T}\log\frac{p(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}{q_{\theta}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}$ is the accumulation of $r_{t^{\prime}}=\log\frac{p(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}{q_{\theta}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}$ that measures the quality of each step’s generation. Intuitively, the generated texts are supposed to have high probabilities under the teacher distribution by increasing $p(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})$, but simultaneously stay diverse by lowering $q_{\theta}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})$. The expectation in Eq. 2 is computed by Monte-Carlo sampling. Full derivation can be found in Appendix A.2. However, policy gradient suffers from high variance and reward hacking [reward_hacking], despite some subsequent solutions [ppo]. Besides, we notice that $R_{t}$ favors short sentences, which causes the student model to output empty responses. Therefore, we propose three strategies to mitigate these problems.
@@
[kd_policy_kd] has found that the single-step generation quality $r_{t}$ is critical to the training variance because the error in the front tokens accumulates along the whole sentence. To pay more attention to $r_{t}$, we re-write $\nabla\mathcal{L}(\theta)$ to decompose $r_{t}$ from $R_{t}$ and directly compute the gradient of $\mathop{\mathbb{E}}_{y_{t}\sim q_{\theta}(t)}[r_{t}]$ (see Appendix A.3 for the full derivation):
@@
We observe reward hacking [reward_hacking] when training with Eq. 2 because $q_{\theta}$ sometimes produces degenerated sentences $\bm{y}$ that receive high scores from the teacher (e.g., repeated phrases) during sampling, especially for small student models. To create a better sampling distribution, we mix the teacher and the student distribution at each time step:
@@
where $\alpha$ controls the strength of the teacher mix-in. Sampling from $\widetilde{p}$ suppresses low-quality generation with the teacher’s help and alleviates reward hacking. We re-write $(\nabla\mathcal{L})_{\text{Single}}$ and $(\nabla\mathcal{L})_{\text{Long}}$ with importance sampling to get to an unbiased estimator of the gradient [off_policy]:
@@
where $w_{t}=\prod_{t^{\prime}=1}^{t}\frac{q_{\theta}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}{\widetilde{p}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}$ is the importance weight. However, $w_{t}$ brings high variance in practice because it requires multiplying per-token importance weight over multiple time steps, and thus the variance of each step accumulates. Therefore, we approximately set $w_{t}\approx\frac{q_{\theta}(y_{t}|\bm{y}_{<t},\bm{x})}{\widetilde{p}(y_{t}|\bm{y}_{<t},\bm{x})}$ to reduce the variance of the estimator in Eq. 5 [deep_rl_chat, offline_rl].
@@ 2.3 Training Algorithm
We start from a student model pre-trained on a large long-document corpus $\mathcal{D}_{\text{PT}}$. Algorithm 1 trains MiniLLM by adapting the student model to a text generation task with dataset $\mathcal{D}$ and supervision from the teacher model, such as an LLM fine-tuned on $\mathcal{D}$ [alpaca, vicuna] or that with good task-generalization [flan-t5, gpt4]. In the training algorithm, we first fine-tune the student model on $\mathcal{D}$ and pick the checkpoint with the lowest loss as an initialization for the following training. Then, we compute the gradients $(\nabla\mathcal{L})_{\text{Single}}$ and $(\nabla\mathcal{L})_{\text{Long}}^{\text{Norm}}$ based on Eq. 5 and Eq. 6, with a clipping strategy [ppo] added to further improve stability. Same as [instruct-gpt], we include a language modeling loss $\mathcal{L}_{\text{PT}}=-\mathop{\mathbb{E}}_{\bm{d}\sim\mathcal{D}_{\text{PT}}}\log q_{\theta}(\bm{d})$ to preserve the model performance on canonical NLP benchmarks. The student model is finally updated using a combination of gradients $(\nabla\mathcal{L})_{\text{Single}}+(\nabla\mathcal{L})_{\text{Long}}^{\text{Norm}}+\nabla\mathcal{L}_{\text{PT}}$. The whole on-policy training pipeline is similar to Reinforcement Learning from Human Feedback (RLHF; instruct-gpt).
@@ 3.1 Experimental Setup
We take instruction-following [instruct-gpt] as the conditional text generation task, where models are trained to generate responses according to the instructions. We fine-tune a large model on the dataset $\mathcal{D}$ consisting of instruction-response pairs as the teacher model. Then, we compare different KD methods on $\mathcal{D}$ by evaluating the student model’s instruction-following performance.
@@
Our student models come from three model families with various sizes: GPT-2 [gpt2] (120M, 340M, 760M), OPT [opt] (1.3B, 2.7B, 6.7B), and LLaMA [llama] (7B). For teacher models of each model family, we use GPT-2-1.5B, OPT-13B, and LLaMA-13B respectively. These models are fine-tuned on $\mathcal{D}$ in advance. We also present the results using GPT-J [gpt-j] as the teacher model in Appendix C.1.
@@
We construct the training data from databricks-dolly-15K [^2] https://github.com/databrickslabs/dolly/tree/master consisting of 15K human-written instruction-response pairs. We filter out samples that exceed the context length of the models. Then, we randomly split 1K and 0.5K samples for validation and testing, respectively, leaving about 12.5K examples for training. For $\mathcal{D}_{\text{PT}}$, we use OpenWebText [openwebtext] for the GPT-2 family and the RoBERTa training corpus [roberta] for other models. We set the teacher-mix-in strength $\alpha=0.2$ throughout the experiments in Eq. 4. We use Rouge-L [rouge] scores on the validation set to search for hyper-parameters because it aligns better with human preference than validation losses [super-natural-instructions]. More details are shown in Appendix B.1.
@@
  • SelfInst [self_inst]: A user-oriented instruction-following set with 252 samples.
  • VicunaEval [vicuna]: The 80 challenging questions used in the Vicuna evaluation.
  • S-NI: The test set of Super-NaturalInstructions [super-natural-instructions] consisting of 9K samples ranging from 119 tasks. Following [ITGPT4], we split the set into 3 subsets whose ground truth response lengths lie in $[0,5]$, $[6,10]$, and $11,+\infty]$. We use the $11,+\infty]$ subset in Section 1 and conduct an analysis on all subsets in Section 3.3.
  • UnNI: We randomly sample 10K samples from the core set of UnnaturalInstructions [uinst] for evaluation. Similar to S-NI, we first conduct the evaluations on the $11,+\infty]$ subset, followed by an analysis of the performance on all subsets in Appendix C.3.
@@
  • R-L: The Rouge-L [rouge] score to measure the precision of the model generation. [super-natural-instructions] has shown that Rouge-L is suitable for large-scale instruction-following evaluation.
  • GPT4: The GPT-4 feedback [mtbench] by asking GPT-4 to compare model-generated responses with the ground truth answers [^3] We use the ChatGPT’s generation [chatgpt] for VicunaEval’s ground truth responses. and raise 1-10 scores for both responses (see Appendix B.2 for the prompt we use). We report the ratio of the total score of model responses and ground truth answers. This metric is only applied to DollyEval, SelfInst, and VicunaEval.
  • Human Evaluation: We conduct human evaluations on the SelfInst dataset following  [ITGPT4] by asking volunteers to compare two responses produced by different models and annotate “Win”, “Tie”, or “Loss”. More human evaluation details can be found in Appendix B.3.
@@
  • KD [distilbert, lightpaff] fine-tunes the student model on $\mathcal{D}$ using the teacher distribution as the supervision at each token step, also known as word-level KD.
  • SeqKD [skd, vicuna, alpaca, ITGPT4, lima] fine-tunes the student model on the data generated by the teacher model.

### v6

# MiniLLM: On-Policy Distillation of Large Language Models

@@ Abstract
Figure 1: The comparison of MiniLLM with the sequence-level KD (SeqKD; 36, 66, 13, 53, 22, 84) in terms of the average GPT-4 feedback score on our evaluation sets. Left: GPT-2-1.5B as the teacher model and GPT-2 125M, 340M, 760M as the student models. Middle: GPT-J 6B as the teacher model and GPT-2 760M, 1.5B, GPT-Neo 2.7B as the student models. Right: OPT 13B as the teacher and OPT 1.3B, 2.7B, 6.7B as the student models.
@@ 1 Introduction
With the rapid development of large language models (LLMs; 9, 24, 8, 14, 50), a common technique to reduce their high computational resource demand is knowledge distillation (KD; 26), where we train a small student model with supervision from a large teacher model. Two categories of KD are commonly applied: black-box KD, where only the teacher-generated texts are accessible, and white-box KD, where the teacher model’s output distribution or intermediate hidden states are also available [33]. Recently, black-box KD has shown promising results in fine-tuning small models on the prompt-response pairs generated by LLM APIs [66, 13, 79, 53]. With the emergence of more open-source LLMs [82, 69], white-box KD becomes more valuable for both research communities and industry sectors because student models receive better signals from the output distribution and hidden states of teacher models, thereby potentially resulting in higher performance. However, white-box KD approaches are mostly studied for small (<< 1B parameters) language understanding models [57, 72], while white-box KD for LLMs is yet to be explored.
In this work, we investigate white-box KD of LLMs where the output distribution of the teacher model is available. We argue that the standard KD objectives [36, 63, 13, 66] are sub-optimal for LLMs that perform tasks in a generative manner. Given the teacher distribution $p(\bm{y}|\bm{x})$ and the student distribution $q_{\theta}(\bm{y}|\bm{x})$ parameterized by $\theta$, standard KD objectives (including several variants for sequence-level models) essentially minimize the approximated forward Kullback-Leibler divergence (KLD) between the teacher and the student distribution, termed as $\operatorname{KL}[p||q_{\theta}]$, which forces $q_{\theta}$ to cover all modes of $p$. For text classification tasks, $\operatorname{KL}[p||q_{\theta}]$ works well because the output space usually consists of a finite number of classes such that both $p(\bm{y}|\bm{x})$ and $q_{\theta}(\bm{y}|\bm{x})$ have few modes. However, for open-ended text generation tasks, which is usually the case of LLM applications, the output spaces are much more complex and $p(\bm{y}|\bm{x})$ can contain many more modes than what $q_{\theta}(\bm{y}|\bm{x})$ can express due to the limited model capacity. Minimizing forward KLD causes $q_{\theta}$ to assign unreasonably high probabilities to the void regions of $p$ [44] and produces very unlikely samples under $p$ during free-run generation [29].
@@
To alleviate this problem, we propose to minimize reverse KLD, $\operatorname{KL}[q_{\theta}||p]$, widely used in computer vision [37] and reinforcement learning [18]. Compared to $\operatorname{KL}[p||q_{\theta}]$, minimizing $\operatorname{KL}[q_{\theta}||p]$ causes $q_{\theta}$ to seek the major modes of $p$, and assign low probabilities to $p$’s void regions [45], as illustrated in Figure 2 and discussed in Section 2.1. In text generation, this means that the student avoids learning too many long-tail variants [27] in the teacher’s distribution to focuses on the generation correctness, which is critical in practical scenarios that require truthfulness and reliability [31]. To optimize $\min_{\theta}\operatorname{KL}[q_{\theta}||p]$, as shown in Section 2.2, we derive its gradient with Policy Gradient [65] and adopt an on-policy training approach. To further stabilize and accelerate training, we propose (1) single-step decomposition to reduce variance, (2) teacher-mixed sampling to alleviate reward hacking, and (3) length normalization to eliminate the length bias. Finally, we introduce the overall KD algorithm in Section 2.3. Our student models are named MiniLLM, indicating our method is suitable for compressing large (generative) language models.
We apply our method to various generative language models [55, 82, 69] with sizes ranging from 120M to 13B in the instruction-following setting [58, 75] that covers a large range of NLP tasks. We use 5 datasets with Rouge-L [42], the GPT-4 feedback, and human judgment for evaluation. Experiments show that MiniLLM consistently outperforms standard KD baselines on all the datasets and scales up well from 120M to 13B models (see Figure 1). More analysis shows that MiniLLM yields lower exposure bias, better calibration, and higher long response generation performance, with neglectable loss of diversity.
@@ 2 Method
We consider conditional text generation where the model produces a response $\bm{y}=\{y_{t}\}_{t=1}^{T}$ conditioning on a prompt $\bm{x}$ sampled from the distribution $p_{\bm{x}}$, which is typically how LLMs perform tasks. We formulate KD as an optimization problem to minimize the difference between a fixed teacher model distribution $p(\bm{y}|\bm{x})$ and a student model distribution $q_{\theta}(\bm{y}|\bm{x})$ parameterized by $\theta$. The standard KD methods approximately [^1] We say “approximately” because for word-level KD, $\bm{y}$ is sampled from the real distribution, not the teacher distribution. For a strong enough teacher model, we can consider the two distributions approximately the same. minimize the forward KLD: $\operatorname{KL}[p||q_{\theta}]=\mathop{\mathbb{E}}_{\bm{x}\sim p_{\bm{x}},\bm{y}\sim p^{\prime}}\log\frac{p(\bm{y}|\bm{x})}{q_{\theta}(\bm{y}|\bm{x})}$, where $p^{\prime}$ can be real data distribution (word-level KD) or teacher distribution $p$ (sequence-level KD). Though widely used, $\operatorname{KL}[p||q_{\theta}]$ tends to overestimate the void regions of $p$ in text generation tasks when $q_{\theta}$ is insufficiently expressive [30]. KD for LLMs fits the case because LLMs perform tasks in a generative manner, such that the low-capacity student models cannot perfectly imitate the complex text generation distribution of the teacher models or humans.
@@ 2.1 MiniLLM: Knowledge Distillation with Reverse KLD
Minimizing reverse KLD has been shown to cause the mode-seeking behavior in generative modeling [29, 48, 12, 37], where $q_{\theta}$ assigns high probabilities to $p$’s large modes and ignore the small ones (illustrated in a toy experiment in Figure 2). In this work, we first study this property for KD of LLMs in text generation. Minimizing forward KLD causes $q_{\theta}$ to place large probability masses on the zero-probability regions of $p$, corresponding to the generation of low-quality text in practice, while reverse KLD focuses on $p$’s major modes, which is crucial to ensure the correctness and faithfulness of text generation. As illustrated in Figure 3, unlike sequence-level KD that minimizes forward KLD [36, 66], MiniLLM that minimizes reverse KLD does not force $q_{\theta}$ to fit all $\bm{y}$ sampled from the teacher distribution $p$. Instead, it encourages the student to generate samples preferred by the teacher within its own capacities, which is more possible to achieve. Interestingly, we also find another perspective of understanding MiniLLM motivated by Inverse Reinforcement Learning [85]. We present the related derivation in Appendix A.1.
@@ 2.2 On-Policy Distillation
We notice that the gradient of the objective function $\mathcal{L}(\theta)$ in Equation (1) can be derived using the Policy Gradient Theorem [78, 23] for on-policy optimization:
@@
where $T=|\bm{y}|$ and $R_{t}=\sum_{t^{\prime}=t}^{T}\log\frac{p(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}{q_{\theta}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}$ is the accumulation of $r_{t^{\prime}}=\log\frac{p(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}{q_{\theta}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}$ that measures the quality of each step’s generation. Intuitively, the generated texts are supposed to have high probabilities under the teacher distribution by increasing $p(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})$, but simultaneously stay diverse by lowering $q_{\theta}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})$. The expectation in Eq. 2 is computed by Monte-Carlo sampling. Full derivation can be found in Appendix A.2. However, policy gradient suffers from high variance and reward hacking [61], despite some subsequent solutions [59]. Besides, we notice that $R_{t}$ favors short sentences, which causes the student model to output empty responses. Therefore, we propose three strategies to mitigate these problems.
@@
[18] has found that the single-step generation quality $r_{t}$ is critical to the training variance because the error in the front tokens accumulates along the whole sentence. To pay more attention to $r_{t}$, we re-write $\nabla\mathcal{L}(\theta)$ to decompose $r_{t}$ from $R_{t}$ and directly compute the gradient of $\mathop{\mathbb{E}}_{y_{t}\sim q_{\theta}(t)}[r_{t}]$ (see Appendix A.3 for the full derivation):
@@
We observe reward hacking [61] when training with Eq. 2 because $q_{\theta}$ sometimes produces degenerated sentences $\bm{y}$ that receive high scores from the teacher (e.g., repeated phrases) during sampling, especially for small student models. To create a better sampling distribution, we mix the teacher and the student distribution at each time step:
@@
where $\alpha$ controls the strength of the teacher mix-in. Sampling from $\widetilde{p}$ suppresses low-quality generation with the teacher’s help and alleviates reward hacking. We re-write $(\nabla\mathcal{L})_{\text{Single}}$ and $(\nabla\mathcal{L})_{\text{Long}}$ with importance sampling to get to an unbiased estimator of the gradient [54]:
@@
where $w_{t}=\prod_{t^{\prime}=1}^{t}\frac{q_{\theta}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}{\widetilde{p}(y_{t^{\prime}}|\bm{y}_{<t^{\prime}},\bm{x})}$ is the importance weight. However, $w_{t}$ brings high variance in practice because it requires multiplying per-token importance weight over multiple time steps, and thus the variance of each step accumulates. Therefore, we approximately set $w_{t}\approx\frac{q_{\theta}(y_{t}|\bm{y}_{<t},\bm{x})}{\widetilde{p}(y_{t}|\bm{y}_{<t},\bm{x})}$ to reduce the variance of the estimator in Eq. 5 [60, 38].
@@ 2.3 Training Algorithm
We start from a student model pre-trained on a large long-document corpus $\mathcal{D}_{\text{PT}}$. Algorithm 1 trains MiniLLM by adapting the student model to a text generation task with dataset $\mathcal{D}$ and supervision from the teacher model, such as an LLM fine-tuned on $\mathcal{D}$ [66, 13] or that with good task-generalization [15, 50]. In the training algorithm, we first fine-tune the student model on $\mathcal{D}$ and pick the checkpoint with the lowest loss as an initialization for the following training. Then, we compute the gradients $(\nabla\mathcal{L})_{\text{Single}}$ and $(\nabla\mathcal{L})_{\text{Long}}^{\text{Norm}}$ based on Eq. 5 and Eq. 6, with a clipping strategy [59] added to further improve stability. Same as [51], we include a language modeling loss $\mathcal{L}_{\text{PT}}=-\mathop{\mathbb{E}}_{\bm{d}\sim\mathcal{D}_{\text{PT}}}\log q_{\theta}(\bm{d})$ to preserve the model performance on canonical NLP benchmarks. The student model is finally updated using a combination of gradients $(\nabla\mathcal{L})_{\text{Single}}+(\nabla\mathcal{L})_{\text{Long}}^{\text{Norm}}+\nabla\mathcal{L}_{\text{PT}}$. The whole on-policy training pipeline is similar to Reinforcement Learning from Human Feedback (RLHF; 51).
@@ 3.1 Experimental Setup
We take instruction-following [51] as the conditional text generation task, where models are trained to generate responses according to the instructions. We fine-tune a large model on the dataset $\mathcal{D}$ consisting of instruction-response pairs as the teacher model. Then, we compare different KD methods on $\mathcal{D}$ by evaluating the student model’s instruction-following performance.
@@
Our student models come from three model families with various sizes: GPT-2 [55] (120M, 340M, 760M), OPT [82] (1.3B, 2.7B, 6.7B), and LLaMA [69] (7B). For teacher models of each model family, we use GPT-2-1.5B, OPT-13B, and LLaMA-13B respectively. These models are fine-tuned on $\mathcal{D}$ in advance. We also present the results using GPT-J [70] as the teacher model in Appendix C.1.
@@
We construct the training data from databricks-dolly-15K [^2] https://github.com/databrickslabs/dolly/tree/master consisting of 15K human-written instruction-response pairs. We filter out samples that exceed the context length of the models. Then, we randomly split 1K and 0.5K samples for validation and testing, respectively, leaving about 12.5K examples for training. For $\mathcal{D}_{\text{PT}}$, we use OpenWebText [20] for the GPT-2 family and the RoBERTa training corpus [43] for other models. We set the teacher-mix-in strength $\alpha=0.2$ throughout the experiments in Eq. 4. We use Rouge-L [42] scores on the validation set to search for hyper-parameters because it aligns better with human preference than validation losses [74]. More details are shown in Appendix B.1.
@@
  • SelfInst [73]: A user-oriented instruction-following set with 252 samples.
  • VicunaEval [13]: The 80 challenging questions used in the Vicuna evaluation.
  • S-NI: The test set of Super-NaturalInstructions [74] consisting of 9K samples ranging from 119 tasks. Following [53], we split the set into 3 subsets whose ground truth response lengths lie in $[0,5]$, $[6,10]$, and $11,+\infty]$. We use the $11,+\infty]$ subset in Section 1 and conduct an analysis on all subsets in Section 3.3.
  • UnNI: We randomly sample 10K samples from the core set of UnnaturalInstructions [28] for evaluation. Similar to S-NI, we first conduct the evaluations on the $11,+\infty]$ subset, followed by an analysis of the performance on all subsets in Appendix C.3.
@@
  • R-L: The Rouge-L [42] score to measure the precision of the model generation. [74] has shown that Rouge-L is suitable for large-scale instruction-following evaluation.
  • GPT4: The GPT-4 feedback [83] by asking GPT-4 to compare model-generated responses with the ground truth answers [^3] We use the ChatGPT’s generation [49] for VicunaEval’s ground truth responses. and raise 1-10 scores for both responses (see Appendix B.2 for the prompt we use). We report the ratio of the total score of model responses and ground truth answers. This metric is only applied to DollyEval, SelfInst, and VicunaEval.
  • Human Evaluation: We conduct human evaluations on the SelfInst dataset following  [53] by asking volunteers to compare two responses produced by different models and annotate “Win”, “Tie”, or “Loss”. More human evaluation details can be found in Appendix B.3.
@@
  • KD [57, 63] fine-tunes the student model on $\mathcal{D}$ using the teacher distribution as the supervision at each token step, also known as word-level KD.
  • SeqKD [36, 13, 66, 53, 84] fine-tunes the student model on the data generated by the teacher model.